OpenStudy (joftheworld):

Ahmed is taking orders for lunch. A slice of pizza cost $3 and a chicken sandwich cost $2. He collects $24 for the group of 10. let x= number of pizza orders. Let y= number of chicken orders. Assuming each person only orders one item. how many people order the chicken sandwich?

OpenStudy (anonymous):

3y+2x=24 y+x=10 x=10-y 3y+2(10-y)= 24 3y+20-2y= 24 3y-2y=4 y=4 4 people order the chicken sandwhcih